Gucchi:

Solve for x: two fifths(x − 2) = 4x. x = two ninths x = −2 x = negative two ninths x = negative nine halves

Vocaloid:

Old question, but just in case you still need help (2/5) * (x-2) = 4x multiply both sides by 5/2 to eliminate 2/5 from the left side (x-2) = (5/2)4x simplifying: x-2 = 10x should be straightforward from there